Camptown Togs, Inc., a children’s clothing man-ufacturer, has always found payroll processing to be costly because it must be done by a clerk. The number of piece–goods coupons received by each employee is collected and the types of tasks performed by each employee are calculated. Not long ago, an industrial engineer designed a system that partially automates the process by means of a scanner that reads the piece–goods coupons. Management is enthusiastic about this system, because it utilizes some personal computer systems that were purchased recently. It is expected that this new automated system will save $45,000 per year in labor. The new system will cost about $30,000 to build and test prior to operation. It is expected that operating costs, including income taxes, will be about $5,000 per year. The system will have a five-year useful life. The expected net salvage value of the system is estimated to be $3,000.

(a) How long does it take to recover the investment?

(b) If the firm’s interest rate is 15% after taxes, what would be the discounted payback period for this project?
Answer #1

A.Initial investment = 30000 + 5000

Saving per year = 45000

Payback period = 0.74 years

B.Interest rate = 15%

Initial investment = 35000

Saving per year = 45000/1.15 = 39130.43